what is the answer to this equation?
3 /(x – 3) + 5 /(x + 4) = (5x + 6)/(x^2 + x - 12)

Answer by checkley77(12844)

3/(x–3)+5/(x+4)=(5x+6)/(x^2+x-12)
[3(x+4)+5(x-3)]/(x-3)(x+4)=(5x+6)/(x^2+x-12)
[3x+12+5x-15]/(x^2+x-12)=(5x+6)/(x^2+x-12) cancel out the denominators.
8x-3=5x+6
8x-5x=6+3
3x=9
x=9/3
x=3 ans.